# Finite–slope universal Rankin–Selberg $p$-adic $L$-function

### v1

@@ 2 Proof and Results
Then the main $R=\mathbb{T}$ theorem of Böckle–Emerton in this context (as formulated in [13, Thm. 3.4], corresponding to [13, Thm. 2.4]) shows that:
@@
    where $T(\bar{\rho}_{2})$ is the localisation at the maximal ideal corresponding to $\bar{\rho}_{2}$ of the prime-to-$p$ Hecke algebra acting on the space $S(1,\mathcal{O})$ of cuspidal pp–adic modular forms of tame level $1$.
By definition of the universal deformation, there is a free rank–$2$ $R(\bar{\rho}_{2})$–module $V_{2}$ equipped with a continuous $G_{\mathbb{Q},\{p\}}$–action lifting $\bar{\rho}_{2}$; we regard this as the universal pp–adic Galois representation of type $\bar{\rho}_{2}$.
We henceforth abbreviate
$R_{2}:=R(\bar{\rho}_{2}),\qquad V_{2}\text{ as above.}$
(1.3) The half–ordinary Rankin–Selberg Panchishkin family.
@@
Since $R_{1}$ and $R_{2}$ are flat complete Noetherian local $\mathcal{O}$–algebras of relative dimensions $1$ and $3$ respectively, their completed tensor product $R$ is again a flat complete Noetherian local $\mathcal{O}$–algebra of relative dimension
@@
On $R$ we have the rank–$4$ RR–module
@@
with its diagonal $G_{\mathbb{Q}}$–action, and the rank–$2$ RR–submodule
@@
which is stable under $G_{\mathbb{Q}_{p}}$ because both factors are. This gives a $G_{\mathbb{Q}_{p}}$–stable filtration
@@
with $V^{+}$ of rank $2$ and $V^{-}$ of rank $2$ over $R$.
This is exactly the datum considered in [13, Ex. 3.17], specialised to the case where the first factor is already ordinary, so that no additional twist is needed to make the ordinary filtration at $p$ appear. In the notation of [13, Ex. 3.17], we are taking
$R=R^{\mathrm{ord}}(\bar{\rho}_{1})\,\widehat{\otimes}_{\mathcal{O}}\,R(\bar{\rho}_{2}),\qquad V=V^{\mathrm{ord}}_{1}\,\widehat{\otimes}_{\mathcal{O}}V_{2},\qquad V^{+}=V^{\mathrm{ord},+}_{1}\,\widehat{\otimes}_{\mathcal{O}}V_{2}.$
@@
Moreover, the interpolation set
@@
is Zariski-dense in $\operatorname{Spf}(R)^{\operatorname{rig}}$ (again [13, Ex. 3.17]). Thus the triple $(R,V,V^{+})$ provides precisely the global deformation space and Panchishkin data required in the hypotheses of Theorem 3.
(1.4) Relation with our chosen neighbourhood $U$.
We now explain how $(R,V,V^{+})$ is related to the eigenvariety $\mathscr{E}$ introduced in the introduction. By construction of the rigid spaces above, the ordinary deformation space $\operatorname{Spf}(R_{1})^{\operatorname{rig}}$ dominates the ordinary locus of the Coleman–Mazur eigencurve (and more generally the parabolic eigenvariety) attached to $\bar{\rho}_{1}$; this is made precise by the $R=\mathbb{T}$ identification between $R_{1}$ and the localised ordinary Hecke algebra, together with the construction of the ordinary eigenvariety (cf. [13, §3.2, Prop. 3.14], and [15, §2] for the Rankin–Selberg set–up). On the other hand, $\operatorname{Spf}(R_{2})^{\operatorname{rig}}$ is, by definition, the universal deformation space for $\bar{\rho}_{2}$ [13, §3.1].
Hence the product
$X:=\operatorname{Spf}(R)^{\operatorname{rig}}\;=\;\operatorname{Spf}(R_{1})^{\operatorname{rig}}\times\operatorname{Spf}(R_{2})^{\operatorname{rig}}$
maps naturally to our Rankin–Selberg eigenvariety $\mathscr{E}$. The general formalism of parabolic eigenvarieties developed by Barrera Salazar–Williams [9] (see in particular [9, Thm. 5.4]), combined with the Rankin–Selberg construction of [15, §2], shows that this map is locally an isomorphism: there exist affinoid neighbourhoods
$U_{R}\subset X,\qquad U_{E}\subset\mathscr{E}$
of the point corresponding to $(f_{0},g_{0})$ such that $U_{R}\cong U_{E}$ as rigid analytic spaces, and $U_{E}$ is identified with the local piece of the parabolic eigenvariety cut out by the Hecke algebra acting on Rankin–Selberg parahoric overconvergent cohomology (cf. [9, Thm. 5.4], together with [13, Prop. 3.14] and [15, §2]).
Shrinking if necessary, we may (and do) choose the affinoid neighbourhood $U\subset\mathscr{E}$ of our fixed classical point $x_{0}=(f_{0},g_{0})$ in Lemma 1 so that
$U\;\simeq\;U_{R}\;\subset\;X=\operatorname{Spf}(R)^{\operatorname{rig}},$
and we henceforth identify $U$ with this connected affinoid subdomain of $\operatorname{Spf}(R)^{\operatorname{rig}}$.
In particular, every classical point $x\in U$ corresponds to a specialisation $(V_{x},V_{x}^{+})$ of the Panchishkin family $(V,V^{+})$ constructed above, and this specialisation coincides with the Rankin–Selberg Galois representation $V(f_{x})\otimes V(g_{x})$ together with the usual half–ordinary local subspace at $p$ (ordinary on the ff–factor, no local condition on the gg–factor), as in [13, Ex. 3.17] and [15, §2]. Thus, after restricting to $U$, the global deformation space and Panchishkin datum appearing in Theorem 3 are entirely furnished by this half–ordinary Rankin–Selberg example of Loeffler.
Step 2. Global triangulation and perfect cohomology complexes
@@
In this step we extract the local $(\varphi,\Gamma)$–module attached to the Panchishkin family $(R,V,V^{+})$ near our fixed point $x_{0}$, and we record the finiteness and base–change properties of its $(\varphi,\Gamma)$–cohomology, following [20, 19].
2.1. The relative $(\varphi,\Gamma)$–module
Let $X:=\operatorname{Spf}(R)^{\operatorname{rig}}$ be the rigid analytic space attached to $R$, and let $V$ be the 44–dimensional RR–linear representation of $G_{\mathbb{Q}}$ constructed in Step 1. As in [19, §2], one associates to the restriction $V|_{G_{\mathbb{Q}_{p}}}$ a family of $(\varphi,\Gamma)$–modules
$\mathscr{D}\;:=\;D_{\operatorname{rig}}^{\dagger}\bigl(V|_{G_{\mathbb{Q}_{p}}}\bigr),$
which is a locally free module of rank $4$ over the relative Robba ring $\mathcal{R}_{X}$, endowed with commuting semilinear actions of $\varphi$ and $\Gamma:=\operatorname{Gal}(\mathbb{Q}_{p}(\mu_{p^{\infty}})/\mathbb{Q}_{p})$. For any point $x\in X$, we write $\mathscr{D}_{x}$ for the fibre; by construction $\mathscr{D}_{x}\cong D_{\operatorname{rig}}^{\dagger}(V_{x})$ as a $(\varphi,\Gamma)$–module over the usual Robba ring over $\mathbb{Q}_{p}$.
Remark 2.
@@ 2.1. The relative $(\varphi,\Gamma)$–module
The existence, functoriality and base–change properties of $D_{\operatorname{rig}}^{\dagger}$ for families of pp–adic Galois representations are given by [19, Thm. 2.2.17]. Concretely, if $A$ is a $\mathbb{Q}_{p}$–affinoid algebra and $V_{A}$ is a finite projective AA–module with a continuous AA–linear action of $G_{\mathbb{Q}_{p}}$, then there is an associated $(\varphi,\Gamma)$–module $D_{\operatorname{rig}}^{\dagger}(V_{A})$ over the relative Robba ring $\mathcal{R}_{A}(\pi)$ (denoted $\mathbf{D}_{\mathrm{rig}}(V_{A})$ in loc. cit.), and for any morphism $A\to B$ of $\mathbb{Q}_{p}$–affinoid algebras one has a natural base–change isomorphism
@@
2.2. Global triangulation à la Liu
By (H1) and (H3), the pair $(V,V^{+})$ defines an rr–Panchishkin family over $R$ in the sense of [13, Def. 2.1, Rem. 2.2(i), Def. 2.3]: at classical points in the interpolation region of loc. cit., the specialisations $V_{x}$ satisfy the rr–Panchishkin condition and $V_{x}^{+}$ is the corresponding Panchishkin subrepresentation. In particular, the Hodge–Tate weight data at $p$ coming from $(V,V^{+})$ endows the family of $(\varphi,\Gamma)$–modules $\mathscr{D}$ with the additional structure making it a weakly refined family in the sense of [20, Def. 1.5] (= Definition 13.1 in loc. cit.). We may therefore apply Liu’s global triangulation theorem.
Theorem 4 (Liu, global triangulation).
Let $X$ be a reduced separated rigid analytic space over $\mathbb{Q}_{p}$, and let $V_{X}$ be a weakly refined family of pp–adic representations of $G_{\mathbb{Q}_{p}}$ over $X$ in the sense of [20, Def. 1.5]. Let $\mathscr{D}_{X}:=D_{\operatorname{rig}}^{\dagger}(V_{X})$ be the associated family of $(\varphi,\Gamma)$–modules. Then there exists a Zariski open and dense subspace $X^{\mathrm{tri}}\subset X$ such that $\mathscr{D}_{X}|_{X^{\mathrm{tri}}}$ admits a global triangulation: that is, a filtration by $(\varphi,\Gamma)$–submodules
@@ 2.2. Global triangulation à la Liu
whose graded pieces $\operatorname{gr}^{i}\mathscr{D}_{X}:=\operatorname{Fil}^{i}/\operatorname{Fil}^{i-1}$ are rank–$1$ $(\varphi,\Gamma)$–modules with prescribed parameters. Moreover, the locus of global triangulation contains all regular non–critical points of $X$ [20, Thm. 1.8] (corresponding to Theorem 13.2 of loc. cit.).
Applying Theorem 4 to $V_{X}=V|_{G_{\mathbb{Q}_{p}}}$ and shrinking $X$ around $x_{0}$ if necessary, we may and do assume that $x_{0}\in X^{\mathrm{tri}}$ and that the Panchishkin condition at $x_{0}$ is non–critical in the sense of [20, Def. 5.29]. By hypotheses (H2) and (H3) we assume that $x_{0}$ is crystalline with small slope and that $V_{x_{0}}^{+}$ defines a non–critical Panchishkin subspace (i.e. the Hodge–Tate weights and Frobenius slopes satisfy the inequalities of [13, Def. 2.1]), so that $x_{0}$ is a regular non–critical point in the sense of [20, Def. 5.29]. Thus there exists a Zariski open neighbourhood
$U^{\mathrm{tri}}\subset X^{\mathrm{tri}}$
of $x_{0}$ and a filtration
@@
of $\mathscr{D}|_{U^{\mathrm{tri}}}$ by $(\varphi,\Gamma)$–submodules, such that each $\operatorname{gr}^{i}\mathscr{D}$ is locally free of rank $1$ over the relative Robba ring over $U^{\mathrm{tri}}$.
Replacing our previously fixed affinoid neighbourhood $U\subset X=\operatorname{Spf}(R)^{\operatorname{rig}}$ (from Lemma 1 and Step 1.4) by a smaller affinoid subdomain if necessary, we may and do assume that
$U\subset U^{\mathrm{tri}}$
and that the above filtration is defined over $U$. We continue to denote this neighbourhood by $U$.
Definition 1.
Over the neighbourhood $U\subset X^{\mathrm{tri}}$, we define
@@
Proposition 1 (Compatibility with the Panchishkin submodule).
For every classical point $x\in U$ corresponding to a pair $(f_{x},g_{x})$, the fibre of $\mathscr{D}^{+}$ at $x$ is canonically identified with the Panchishkin submodule of $D_{\operatorname{rig}}^{\dagger}(V_{x})$:
$(\mathscr{D}^{+})_{x}\;\cong\;D_{\operatorname{rig}}^{\dagger}(V_{x}^{+})\subset D_{\operatorname{rig}}^{\dagger}(V_{x}).$
@@
By the construction of the rr–Panchishkin family $(R,V,V^{+})$ in [13, §2], the submodule $V^{+}\subset V|_{G_{\mathbb{Q}_{p}}}$ is locally free of rank $r$ over $R$ and, for each classical point $x$ in the interpolation region of loc. cit., the fibre $V_{x}^{+}\subset V_{x}$ is the Panchishkin subrepresentation in the sense of [13, Def. 2.1, Rem. 2.2(i), Def. 2.3]. In particular, $V_{x}^{+}$ is the unique $G_{\mathbb{Q}_{p}}$–stable subspace of $V_{x}$ of dimension $\dim V_{x}^{c=+1}-r$ whose Hodge–Tate weights are strictly larger than those of $V_{x}/V_{x}^{+}$, and the associated $(\varphi,\Gamma)$–module $D_{\operatorname{rig}}^{\dagger}(V_{x}^{+})$ is a saturated submodule of $D_{\operatorname{rig}}^{\dagger}(V_{x})$ of rank $r$.
@@
$(\operatorname{Fil}^{r}\mathscr{D})_{x}\;=\;D_{\operatorname{rig}}^{\dagger}(V_{x}^{+}).$
Since $\mathscr{D}^{+}:=\operatorname{Fil}^{r}\mathscr{D}$ by definition, this proves the claim. ∎
2.3. Finiteness and base change for $(\varphi,\Gamma)$–cohomology
We now record the finiteness and base–change properties of the $(\varphi,\Gamma)$–cohomology and Iwasawa cohomology of $\mathscr{D}$ and $\mathscr{D}^{+}$ over $U$, following [19]. Let us briefly recall the complexes used by Kedlaya–Pottharst–Xiao.
For a $(\varphi,\Gamma)$–module $\mathsf{M}$ over a relative Robba ring $\mathcal{R}_{A}(\pi)$, we denote by $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})$ the complex computing the usual $(\varphi,\Gamma)$–cohomology, and by $C^{\bullet}_{\psi}(\mathsf{M})$ the complex computing Iwasawa cohomology, as defined in [19, §4].
Theorem 5 (Kedlaya–Pottharst–Xiao).
Let $A$ be an affinoid $\mathbb{Q}_{p}$–algebra and let $\mathsf{M}$ be a $(\varphi,\Gamma)$–module over the relative Robba ring $\mathcal{R}_{A}(\pi)$. Then:

### v2

@@ 2.1.2 The unrestricted universal deformation ring and representation for $\bar{\rho}_{2}$
The main $R=\mathbb{T}$ theorem of Böckle–Emerton in this context (as formulated in [13, Thm. 3.4]) shows that:
  1. (a)
@@
    where $T(\bar{\rho}_{2})$ is the localisation at the maximal ideal corresponding to $\bar{\rho}_{2}$ of the prime-to-$p$ Hecke algebra acting on the space $S(1,\mathcal{O})$ of cuspidal $p$-adic modular forms of tame level $1$.
By definition of the universal deformation, there is a free rank two $R(\bar{\rho}_{2})$-module $V_{2}$ equipped with a continuous $G_{\mathbb{Q},\{p\}}$-action lifting $\bar{\rho}_{2}$; we regard this as the universal $p$-adic Galois representation of type $\bar{\rho}_{2}$. We henceforth abbreviate
$R_{2}:=R(\bar{\rho}_{2}).$
2.1.3 The half–ordinary Rankin–Selberg Panchishkin family
@@
Since $R_{1}$ and $R_{2}$ are flat complete Noetherian local $\mathcal{O}$-algebras of relative dimensions $1$ and $3$ respectively, their completed tensor product $R$ is again a flat complete Noetherian local $\mathcal{O}$-algebra of relative dimension
@@
On $R$ we have the rank four $R$-module
@@
with its diagonal $G_{\mathbb{Q}}$-action, and the rank two $R$-submodule
@@
which is stable under $G_{\mathbb{Q}_{p}}$ because both factors are. This gives a $G_{\mathbb{Q}_{p}}$-stable filtration
@@
with $V^{+}$ and $V^{-}$ of rank $2$ over $R$.
This is exactly the datum considered in [13, Ex. 3.17], specialised to the case where the first factor is already ordinary. In the notation of loc. cit., we are taking
$R=R^{\mathrm{ord}}(\bar{\rho}_{1})\widehat{\otimes}_{\mathcal{O}}R(\bar{\rho}_{2}),\qquad V=V_{1}^{\mathrm{ord}}\widehat{\otimes}_{\mathcal{O}}V_{2},\qquad V^{+}=V_{1}^{\mathrm{ord},+}\widehat{\otimes}_{\mathcal{O}}V_{2}.$
Loeffler proves in [13, Ex. 3.17] that $(V,V^{+})$ is a 0-Panchishkin family over $R$ in his sense: for every arithmetic point $\kappa$ of $\operatorname{Spf}(R)^{\operatorname{rig}}$ corresponding to a pair $(f,g)$ of classical eigenforms, the specialisation $(V_{\kappa},V_{\kappa}^{+})$ coincides with the Rankin–Selberg Galois representation $V(f)\otimes V(g)$ together with the usual half–ordinary subspace at $p$. Moreover, the interpolation set
@@
is Zariski-dense in $\operatorname{Spf}(R)^{\operatorname{rig}}$. Thus the triple $(R,V,V^{+})$ provides the global deformation space and Panchishkin data referred to in Theorem 1.4.
2.1.4 Relation with the eigenvariety and with the ordinary locus in $U$
We now explain how the universal half–ordinary Panchishkin family of [13, Ex. 3.17] interacts with the parabolic eigenvariety $\mathscr{E}$ and with the neighbourhood $U$ of our fixed point $x_{0}$.
Let $E$ denote the parabolic eigenvariety of [9] attached to $G=\operatorname{GL}_{2}\times\operatorname{GL}_{2}$, with weight space $\mathcal{W}$ and weight map $w:E\to\mathcal{W}$. We keep the notation
$X_{1}:=\operatorname{Spf}(R_{1})^{\operatorname{rig}},\qquad X_{2}:=\operatorname{Spf}(R_{2})^{\operatorname{rig}},\qquad X:=X_{1}\times X_{2}=\operatorname{Spf}(R)^{\operatorname{rig}}.$
By construction, $X_{1}$ dominates the ordinary locus of the Coleman–Mazur eigencurve (and more generally the parabolic eigenvariety) attached to $\bar{\rho}_{1}$; this is a consequence of Hida theory together with the $R=\mathbb{T}$ theorems used in [13, §3.2]. The space $X_{2}$ is the universal deformation space for $\bar{\rho}_{2}$.
On the locus where the first factor is ordinary at $p$ there is a natural morphism
$X=X_{1}\times X_{2}\longrightarrow E$
whose points correspond to pairs of classical cusp forms $(f,g)$ with $f$ ordinary at $p$, together with an ordinary refinement on the $f$-factor. On this locus the specialisation $(V_{x},V_{x}^{+})$ of $(V,V^{+})$ is the Rankin–Selberg representation $V(f)\otimes V(g)$ together with the usual half–ordinary subspace at $p$, as in [13, Ex. 3.17] and [15].
Let $E^{\mathrm{ord}}\subset E$ denote the ordinary locus for the first factor, and set
$U^{\mathrm{ord}}:=U\cap E^{\mathrm{ord}}.$
This is an admissible open subset of $U$. Since ordinary classical points are Zariski dense in the ordinary eigenvariety and $U$ contains a Zariski-dense set of classical points, we may, after shrinking $U$ slightly if necessary, assume that $U^{\mathrm{ord}}$ is non-empty and that classical points are still Zariski dense in $U^{\mathrm{ord}}$.
On $E^{\mathrm{ord}}$ the eigenvariety machine of [9, 29], together with the Rankin–Selberg construction of [15], shows that the map $X\to E$ above identifies an affinoid neighbourhood of any classical half–ordinary point with an affinoid subdomain of $X$ (see [9, Thm. 5.4], [13, §3.2] and [15, §2]). In particular, there exists an affinoid subdomain
$U_{R}\subset X$
and an isomorphism of rigid spaces over weight space
$\iota:U^{\mathrm{ord}}\xrightarrow{\ \sim\ }U_{R}\subset X$
such that, for every classical point $x\in U^{\mathrm{ord}}$ corresponding to a pair $(f_{x},g_{x})$ with $f_{x}$ ordinary at $p$, the specialisation of the universal family $(V,V^{+})$ at $\iota(x)$ coincides with the Rankin–Selberg representation
$V(f_{x})\otimes V(g_{x})$
together with its standard half–ordinary subspace at $p$.
Two points are important here.
First, the finite-slope point $x_{0}=(f_{0},g_{0})$ does not lie in the image of $X\to E$ when $f_{0}$ is non-ordinary at $p$. Thus $x_{0}$ itself does not belong to $U^{\mathrm{ord}}$, and we do not try to realise $x_{0}$ as a point of the rigid fibre of $R$.
@@ 2.1.4 Relation with the eigenvariety and with the ordinary locus in $U$
[…]
2.2 Step 2: Global triangulation and perfect cohomology complexes
@@
In this step we work entirely over the fixed affinoid neighbourhood $U\subset E$ of $x_{0}$. We recall the family of local Galois representations on $U$, attach to it a family of $(\varphi,\Gamma)$-modules, and record the finiteness and base-change properties of the associated cohomology complexes, following [20, 19].
2.2.1 The relative $(\varphi,\Gamma)$-module
By the eigenvariety machine of Hansen and Barrera Salazar–Williams (see [29, Thm. 4.5.1] and [9, §5]), there exists a locally free rank four $\mathcal{O}(U)$-module $V_{U}$ equipped with a continuous $\mathcal{O}(U)$-linear action of $G_{\mathbb{Q}}$, unramified outside $p$, such that for every classical point $x\in U$ corresponding to a pair of cuspidal eigenforms $(f_{x},g_{x})$ the fibre $V_{x}$ is canonically isomorphic to the Rankin–Selberg tensor
$V(f_{x})\otimes V(g_{x}).$
The $U_{p}$-eigenvalues on $U$ give an analytic function $\alpha:U\to\mathcal{O}(U)^{\times}$ encoding a refinement of $V_{U}\big|_{G_{\mathbb{Q}_{p}}}$; this makes $V_{U}\big|_{G_{\mathbb{Q}_{p}}}$ into a weakly refined family in the sense of [20, Def. 1.5], cf. the discussion in [20, §5] and [29, §4.5].
Let
$\mathscr{D}:=D_{\operatorname{rig}}^{\dagger}\bigl(V_{U}\big|_{G_{\mathbb{Q}_{p}}}\bigr)$
be the associated family of $(\varphi,\Gamma)$-modules over the relative Robba ring $\mathcal{R}_{\mathcal{O}(U)}$; cf. [19, Thm. 2.2.17]. For any point $x\in U$ we write $\mathscr{D}_{x}$ for the fibre; by construction
$\mathscr{D}_{x}\cong D_{\operatorname{rig}}^{\dagger}(V_{x})$
as a $(\varphi,\Gamma)$-module over the usual Robba ring over $\mathbb{Q}_{p}$.
Remark 2.1.
@@ 2.2.1 The relative $(\varphi,\Gamma)$-module
The existence, functoriality and base-change properties of $D_{\operatorname{rig}}^{\dagger}$ for families of $p$-adic Galois representations are given by [19, Thm. 2.2.17]. Concretely, if $A$ is a $\mathbb{Q}_{p}$-affinoid algebra and $V_{A}$ is a finite projective $A$-module with a continuous $A$-linear action of $G_{\mathbb{Q}_{p}}$, then there is an associated $(\varphi,\Gamma)$-module $D_{\operatorname{rig}}^{\dagger}(V_{A})$ over the relative Robba ring $\mathcal{R}_{A}$, and for any morphism $A\to B$ of $\mathbb{Q}_{p}$-affinoid algebras one has a natural base-change isomorphism
@@
2.2.2 Global triangulation à la Liu
We recall Liu’s global triangulation theorem in the form needed here.
Theorem 2.2 (Liu, global triangulation).
Let $X$ be a reduced separated rigid analytic space over $\mathbb{Q}_{p}$, and let $V_{X}$ be a weakly refined family of $p$-adic representations of $G_{\mathbb{Q}_{p}}$ over $X$ in the sense of [20, Def. 1.5]. Let $\mathscr{D}_{X}:=D_{\operatorname{rig}}^{\dagger}(V_{X})$ be the associated family of $(\varphi,\Gamma)$-modules. Then there exists a Zariski open and dense subspace $X^{\mathrm{tri}}\subset X$ such that $\mathscr{D}_{X}\big|_{X^{\mathrm{tri}}}$ admits a global triangulation: that is, a filtration by $(\varphi,\Gamma)$-submodules
@@ 2.2.2 Global triangulation à la Liu
whose graded pieces $\operatorname{gr}^{i}\mathscr{D}_{X}:=\operatorname{Fil}^{i}/\operatorname{Fil}^{i-1}$ are rank one $(\varphi,\Gamma)$-modules with prescribed parameters. Moreover, the locus of global triangulation contains all regular non-critical points of $X$ [20, Thm. 1.8].
We apply Theorem 2.2 with $X=U$ and $V_{X}=V_{U}\big|_{G_{\mathbb{Q}_{p}}}$. By hypotheses (H2) and (H3), the point $x_{0}$ is crystalline with small slope and the local conditions at $p$ define a non-critical Panchishkin situation in the sense of [20, Def. 5.29]; in particular $x_{0}$ is a regular non-critical point. Hence there is a Zariski open neighbourhood $U^{\mathrm{tri}}\subset U$ of $x_{0}$ on which $\mathscr{D}$ admits a global triangulation.
After shrinking $U$ if necessary, we may and do assume that $U\subset U^{\mathrm{tri}}$ and that there exists a filtration
@@
of $\mathscr{D}$ by $(\varphi,\Gamma)$-submodules such that each $\operatorname{gr}^{i}\mathscr{D}$ is locally free of rank one over the relative Robba ring over $U$.
Definition 2.3.
Over the neighbourhood $U$ we define
@@
Proposition 2.4 (Compatibility with the Panchishkin local condition).
[…]
$(\mathscr{D}^{+})_{x}\cong D_{\operatorname{rig}}^{\dagger}(V_{x}^{\mathrm{ord},+}),$
where $V_{x}^{\mathrm{ord},+}\subset V_{x}$ is the rank $r$ ordinary $G_{\mathbb{Q}_{p}}$-subrepresentation coming from the Hida deformation of $\bar{\rho}_{1}$.
@@
[…]
2.2.3 Finiteness and base change for $(\varphi,\Gamma)$-cohomology
We now record the finiteness and base-change properties of the $(\varphi,\Gamma)$-cohomology and Iwasawa cohomology of $\mathscr{D}$ and $\mathscr{D}^{+}$ over $U$, following [19]. Let us briefly recall the complexes used by Kedlaya–Pottharst–Xiao. For a $(\varphi,\Gamma)$-module $\mathsf{M}$ over a relative Robba ring $\mathcal{R}_{A}$, we denote by $C^{\bullet}_{\varphi,\Gamma}(\mathsf{M})$ the complex computing the usual $(\varphi,\Gamma)$-cohomology, and by $C^{\bullet}_{\psi}(\mathsf{M})$ the complex computing Iwasawa cohomology, as defined in [19, §4].
Theorem 2.5 (Kedlaya–Pottharst–Xiao).
Let $A$ be an affinoid $\mathbb{Q}_{p}$-algebra and let $\mathsf{M}$ be a $(\varphi,\Gamma)$-module over the relative Robba ring $\mathcal{R}_{A}$. Then: